OpenStudy (anonymous):

Lim as x --> 0 (2x^2-3x)/2sinx

OpenStudy (anonymous):

Think of it as: \[ \frac{2x-3}{2}\times \frac{x}{\sin x} \]

OpenStudy (anonymous):

Or even as \[ \frac{2x-3}{2}\div \frac{\sin x}{x} \]

OpenStudy (anonymous):

WOW! I think that does it!

OpenStudy (anonymous):

Good luck.

OpenStudy (anonymous):

Sinx/x =1

OpenStudy (anonymous):

2(0)-3/2 = -3/2

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

Much Love! <3